OpenStudy (ellaella02):

What is the slope-intercept form of the function that contains the point (–3, 2) and has a slope of –2? y = x +

OpenStudy (ellaella02):

PLEASE HELP I WILL MEDAL AND FAN AND EVERYTHING

OpenStudy (ellaella02):

@Michele_Laino

OpenStudy (ellaella02):

@Hero

OpenStudy (ellaella02):

@Nnesha

OpenStudy (michele_laino):

here we have to apply this formula: \(y-y_0=m(x-x_0)\) where \(m\) is the slope, \(x_0=-3,y_0=2\)

OpenStudy (sweetburger):

slope intercept is y=mx + b where m is the slope and b is the y intercept therefore you know what m is as it is given in the question. Now plug in the set of given points for x and y and proceed to solve for b.

OpenStudy (ellaella02):

Um how do I do that?

OpenStudy (michele_laino):

it is simple, you have to replace \(m=-2\), \(x_0=-3,y_0=2\) so we get: \(y-2=-2(x+3)\) please simplify

OpenStudy (ellaella02):

Would it be x=−1/2 y−2

OpenStudy (ellaella02):

y=−2x−4

OpenStudy (michele_laino):

hint: here is the next step: \(y-2=-2x-6\) correct!

OpenStudy (ellaella02):

y=−2x−4 is my answer ?

OpenStudy (michele_laino):

correct! You got the right equation!

OpenStudy (ellaella02):

Thank you so much (:

OpenStudy (michele_laino):

:)